귤 organic-2 □ kelin-Bing Organic chemistry (2e)-De bb-Bing āorgo 333 HW1(1).pdf orgo 333 HW1(1).pdf 1+ × ﹀ 0%20%203 33%20HW 1 %20(1).pdf 11. (1 pt) The signal below is a triplet. Determine the coupling constant if the spectrometer frequency is 400 MHz. Show your calculations H Ha 5.75 12. (6 pts) Determine the structures. Characterize as many compounds as you wish from the Workbook Type here to search 6-18 AM 3/2/2018
Answer #1

Here What I got,

Triplet has center pick chemical shift =5.7698

Adjacent picks are both same chemical shift =5.7819

Spectrometer frequency =400 mHz

Coupling Constant (J) =Difference between the Centre pick and any adjacent pick * MHz

= 5.7819-5.7698\times 400=4.48 MHz

hence coupling constant is.4.48 MHz